# Multimodal Data Integration for Oncology in the Era of Deep Neural Networks: A Review

### v2

@@ Abstract
Cancer has relational information residing at varying scales, modalities, and resolutions of the acquired data, such as radiology, pathology, genomics, proteomics, and clinical records. Integrating diverse data types can improve the accuracy and reliability of cancer diagnosis and treatment. Traditional methods typically focus on partial or unimodal information about biological systems at individual scales and fail to encapsulate the complete spectrum of the heterogeneous nature of data. Deep neural networks have facilitated the development of sophisticated multimodal data fusion approaches that can extract and integrate relevant information from multiple sources. Recent deep learning frameworks such as Graph Neural Networks (GNNs) and Transformers have shown remarkable success in multimodal learning. This review article provides an in-depth analysis of the state-of-the-art in GNNs and Transformers for multimodal data fusion in oncology settings, highlighting notable research studies and their findings. We also discuss the foundations of multimodal learning, inherent challenges, and opportunities for integrative learning in oncology. By examining the current state and potential future developments of multimodal data integration in oncology, we aim to demonstrate the promising role that multimodal neural networks can play in cancer prevention, early detection, and treatment through informed oncology practices in personalized settings.
1 Introduction
[…]
Deep learning (DL), a subset of machine learning (ML), uses neural networks with multiple layers to analyze and process large amounts of data. This technique has been used to develop computer-aided diagnostic (CAD) systems that automatically detect and classify cancerous regions in medical images, such as mammograms and computerized tomography (CT) scans. DL has also been used to analyze molecular data to predict a patient’s treatment response and identify new biomarkers for cancer diagnosis. With the growing amount of data generated in oncology, DL is becoming an increasingly important tool for understanding cancer, making diagnoses, predicting clinical outcomes, assessing the effectiveness of different therapies, and developing new treatments [2].
@@ 1 Introduction
Figure 1: Number of publications involving deep learning, graph neural networks (GNNs), GNNs in the medical domain, overall multimodal and multimodal in biomedical and clinical sciences in the period 2014-2023 [3].
1.1 Multimodal Learning (MML)
One particularly promising approach is multimodal learning (MML), which combines information from multiple sources (or modalities) to improve the accuracy and reliability of a given task [4]. MML has gained notable popularity in recent years, as evidenced by the increasing number of publications related to MML (see Fig. 1). With the increasing availability of multimodal data, MML has become an important tool for advancing understanding of complex diseases like cancer. Multimodal data fusion has the potential to enable the integration of diverse data types, such as medical images, molecular data, and electronic health records (EHR), to gain a more comprehensive understanding of a given problem. Combining information from multiple sources makes it possible to extract and integrate relevant features that may be missed when each data modality is analyzed in isolation.
Recent approaches to multimodal data fusion, facilitated by DNNs, can learn from computer vision (CV) and natural language processing (NLP) data simultaneously. Contrastive Language-Image Pre-training (CLIP) and Generative Pre-training Transformer (GPT-4) introduced by OpenAI are multimodal foundation models [5, 6]. Another example of MML is the Foundational Language And Vision Alignment Model (FLAVA) [7], which combined vision and language representation learning for multimodal reasoning. MML has also been applied in oncology. Recently proposed RadGenNets model predicts gene mutations in Non-small cell lung cancer (NSCLC) patients by integrating clinical and genomics data, Positron Emission Tomography (PET) scans, and gene mutation data using a fusion of Convolutional Neural Networks (CNNs) and Dense Neural Networks [8].
1.2 Prior Work
Several survey papers have been published on the topic of multimodal data learning [9, 10, 11, 12], which provide an overview of the current state of the field and have identified key trends and challenges. In oncology, researchers have traditionally pursued multimodal integration through classical ML approaches. Recently, multimodal integration using DNNs such as graph neural networks (GNNs) and Transformers has gained significant traction. GNNs and Transformers have been proposed for various tasks in oncology, including tumor classification [13], prognosis [14], and treatment response assessment [15].
1.3 Contributions
The success of multimodal data fusion models in traditional ML areas, such as CV and NLP, shows their potential for improving cancer care [9]. While many recent surveys cover the topic of MML, there is a lack of surveys specifically focused on using multimodal DL models (such as GNNs and Transformers) in oncology settings. In the following, we provide the contributions of this article:
@@ 1.3 Contributions
    Identifying large-scale MML approaches in oncology. We provide an overview of the current state-of-the-art in using DNNs, specifically GNNs and Transformers, for multimodal data fusion in oncology.
@@
    Highlighting the challenges and limitations of multimodal oncology data fusion. We discuss the challenges and limitations of implementing multimodal data fusion models in oncology, including the need for large datasets, the complexity of integrating diverse data types, data alignment, missing data modalities and samples.
@@
    Providing a taxonomy for describing multimodal architectures. We present a comprehensive taxonomy for describing multimodal architectures, including both traditional and DL methods, to facilitate future research in this area.
@@
    Identifying key contributions and future directions for multimodal data fusion in oncology. Finally, we explore the various ways of integrating multimodal oncology data using neural networks. We identify GNNs and Transformers as a potential solution for across-the-spectrum multimodal integration and present the associated challenges.
2 Fundamentals of Multimodal Learning
Figure 2: From population to single cell, cancer data modalities at scales.
The word modality represents the expression of an entity or a particular form of sensory perception, such as the visual actions of the characters, sounds of the dialogues being spoken, or the background music [16]. A collective notion of these modalities is called multi-modality [11].
@@ 2.1 Data Modalities in Oncology
In oncology, data is inherently multimodal, as it includes information from multiple sources or modalities that are related in various ways. Fig. 2 provides a view of multiple modalities of cancer at various scales, from the population level to the single-cell analysis. Oncology data can be broadly classified into three categories: clinical, molecular, and imaging, where each category provides complementary information about the patient’s disease. Fig. 3 highlights different modalities within clinical, molecular, and imaging categories.
Traditional methods to study cancer are targeted toward individual data modalities (e.g., EHR [17], radiology [18], pathology [19], and molecular, including genomics [20], transcriptomics [21], proteomics [22], etc.). Multi-scale or multi-resolution studies aim to analyze the disease across different scales that may comprise cellular, tissue, and organismic levels [23] or identify the infection outcome of disease across tissues and organs scales [24]. Multimodal analysis endeavors to gain holistic insights into the disease through multiple measurement methods. For the confines of this review, we will collectively call all these multi-categories as multimodal data.
@@ 2.1.1 Molecular Data
Molecular modalities provide information about the underlying genetic changes and alterations in the cancer cells [25]. Efforts toward integrating molecular data resulted in the multi-omics research field. Two main areas of molecular analysis in oncology are proteomics and genomics. Proteomics involves the study of proteins and their changes in response to cancer, and it provides information about the biological processes taking place in cancer cells. Genomics involves the study of the entire genome of cancer cells, including changes in DNA sequence, gene expression, and structural variations [9]. Other molecular modalities include transcriptomics, pathomics, radiomics, and their combinations, radiogenomics and proteogenomics. Many publicly available datasets provide access to molecular data, including the Proteomics Data Commons (PDC) for proteomics data and the Genome Data Commons (GDC) for genetic data [26, 27].
Figure 3: A landscape of oncology data (sub)modalities acquired for cancer care.
@@ 2.1.2 Imaging Data
Imaging modalities play a crucial role in diagnosing and monitoring cancer. The imaging category can be divided into two main categories: radiological and histopathological. Radiological imaging encompasses various techniques such as X-rays, CT scans, and Magnetic resonance imaging (MRI), which provide information about the location and extent of cancer within the body. These images can be used to determine the size and shape of a tumor, monitor its growth, and assess the effectiveness of treatments. Histopathological imaging involves the examination of tissue samples obtained through biopsy or surgery [28]. These images (referred to as WSI) provide detailed information about the micro-structural changes in the cancer cells and can be used to diagnose cancer and determine its subtype.
@@ 2.1.3 Clinical Data
Clinical data provides information about the patient’s medical history, physical examination, and laboratory results. Clinical data can be divided into time-series data, electronic health records (EHR), and clinician notes. Time-series data refers to the collection of clinical data over time, such as repeated blood tests. This data informs the changes in the patient’s condition and monitor the disease progression [29]. EHR consists of digital records of a patient’s health information stored in a centralized database. They provide a comprehensive view of a patient’s medical history, past diagnoses, treatments, and laboratory test results, which help in understanding the disease [30]. Clinical notes are part of EHR and refer to the written records of a patient’s medical history and examinations, as well as the observations and recommendations of the healthcare provider[29].
2.2 Taxonomy of Multimodal Learning
We follow the taxonomy proposed by William et al. [16] (see Fig. 4), which defines five main stages of multimodal classification - preprocessing, feature extraction, data fusion, primary learner, and final classifier, as given below:
@@ 2.2.1 Pre-processing
Pre-processing involves modifying the input data to a suitable format before being fed into the model for training. It includes data cleaning, normalization, class balancing, and augmentation. Data cleaning removes unwanted noise or bias, errors, and missing data points [31]. Normalization scales the input data within a specific range to ensure that each modality contributes equally to the training [32]. Class balancing is done in cases where one class may have a significantly larger number of data points than another, resulting in model bias towards the dominant class. Data augmentation artificially increases the size of the dataset by generating new samples based on the existing data to improve model’s robustness and generalizability [31].
@@ 2.2.2 Feature Extraction
The different modalities may have different features, and extracting relevant features can improve the model’s learning. Several manual and automated feature engineering techniques generate representations (or embeddings) for each data modality. Feature engineering involves designing features relevant to the task and extracting them from the input data. This can be time-consuming but may allow the model to incorporate prior knowledge about the problem. Text encoding techniques, such as bag-of-words, word embeddings, topic models [33, 34], transform textual data into a numerical representation, which can be used as input to an ML model [35]. ML model-generated filters are a feature extraction technique particularly useful for image and video data [36]. These filters are learned automatically by DNNs during the training process and can capture relevant features such as edges, textures, or shapes [36].
@@ 2.2.3 Data Fusion
Data fusion combines raw features, extracted features, or class prediction vectors from multiple modalities to create a single data representation. Fusion enables the model to use the complementary information provided by each modality and improve its learning. Data fusion can be done using early, late, or intermediate fusion. Section 2.3 presents a detailed discussion of these fusion stages. The choice of fusion technique depends on the characteristics of the data and the specific problem being addressed [37].
@@ 2.2.4 Primary Learner
The primary learner stage involves training the model on the pre-processed, feature-extracted data. Depending on the specific problem and data, the primary learner can be implemented using various ML techniques. DNNs are a popular choice for primary learners in MML because they can automatically learn high-level representations from the input data and have demonstrated state-of-the-art performance in many applications. CNNs are often used for image and video data, while recurrent neural networks (RNNs) and Transformers are commonly used for text and sequential data. The primary learner can be implemented independently for each modality or shared between modalities, depending on the specific problem and data.
@@ 2.2.5 Final Classifier
The final stage of MML is the classifier, which produces category labels or class scores and can be trained on the output of the primary learner or the fused data. This stage can include anything from a simple shallow neural network, decision tree, or an ensemble model [16]. Ensemble methods, such as stacking or boosting, are often used to improve the performance of the final classifier. Stacking involves training multiple models and then combining their predictions at the output stage, while boosting involves iteratively training weak learners and adjusting their weights based on the errors made by previous learners [38].
2.3 Data Fusion Stages
@@
Multimodal data fusion can be performed at different levels, including early (feature level), intermediate (model level), and late (decision level), as illustrated in Fig. 4. Each fusion stage has its own advantages and challenges, and the choice of fusion stage depends on the characteristics of the data and the problem at hand. The following sections will explore each fusion stage and discuss its applications and limitations.
@@
Figure 4: Taxonomy, stages, and techniques of multimodal data fusion presented in this survey. Early, late, cross-modality fusion methods integrate individual modality features before, after, or at the primary learning step, respectively.
@@
The early fusion approach involves merging the features extracted from different modalities before the model is trained and combining them into a single feature vector. In this method, the feature vectors of the different modalities are combined into a single vector, which is used as the input to the ML model [16]. This approach can be used when the modalities have complementary information and can be easily aligned, such as combining visual and audio features in a video analysis application. The main challenge with early fusion is ensuring that the features extracted from different modalities are compatible and provide complementary information.

### v3

@@ Abstract
Cancer research encompasses data across various scales, modalities, and resolutions, from screening and diagnostic imaging to digitized histopathology slides to various types of molecular data and clinical records. The integration of these diverse data types for personalized cancer care and predictive modeling holds the promise of enhancing the accuracy and reliability of cancer screening, diagnosis, and treatment. Traditional analytical methods, which often focus on isolated or unimodal information, fall short of capturing the complex and heterogeneous nature of cancer data. The advent of deep neural networks has spurred the development of sophisticated multimodal data fusion techniques capable of extracting and synthesizing information from disparate sources. Among these, Graph Neural Networks (GNNs) and Transformers have emerged as powerful tools for multimodal learning, demonstrating significant success. This review delves into the recent advancements in GNNs and Transformers for the fusion of multimodal data in oncology, spotlighting key studies and their pivotal findings. We explore the foundational principles of multimodal learning and discuss the unique challenges it faces, such as data heterogeneity and integration complexities, alongside the opportunities it presents for a more nuanced and comprehensive understanding of cancer. By surveying the landscape of multimodal data integration in oncology, our goal is to underline the transformative potential of multimodal GNNs and Transformers. These models are poised to revolutionize cancer prevention, early detection, and treatment strategies by fostering informed, personalized oncology practices. By addressing both the technological advancements and the methodological innovations, we aim to chart a course for future research in this promising field. Our review highlights not only the current state of multimodal modeling applications in oncology but also sets the stage for their evolution, encouraging further exploration and development in personalized cancer care.
Cancer represents a significant global health challenge, characterized by the uncontrolled growth of abnormal cells, leading to millions of deaths annually. In 2023, the United States had around 1.9 million new cancer diagnoses, with cancer being the second leading cause of death and anticipated to result in approximately 1670 deaths daily [1]. However, advancements in oncology research hold the promise of preventing nearly 42% of these cases through early detection and lifestyle modifications. The complexity of cancer, involving intricate changes at both the microscopic and macroscopic levels, requires innovative approaches to its understanding and management. In recent years, the application of machine learning (ML) techniques, especially deep learning (DL), has emerged as a transformative force in oncology. DL employs deep neural networks to analyze vast datasets, offering unprecedented insights into cancer’s development and progression [2, 3, 4, 5, 6]. This approach has led to the development of computer-aided diagnostic systems capable of detecting and classifying cancerous tissues in medical images, such as mammograms and MRI scans, with increasing accuracy. Beyond imaging, DL also plays a crucial role in analyzing molecular data, aiding in the prediction of treatment responses, and the identification of new biomarkers [7, 8, 9, 10, 11, 12, 13]. As the volume of oncology data continues to grow, DL stands at the forefront of this field, enhancing our understanding of cancer, improving diagnostic precision, predicting clinical outcomes, and paving the way for innovative treatments. This review explores the latest advancements in DL applications within oncology, highlighting its potential to revolutionize cancer care [14, 15, 16, 17].
@@
Figure 1: Number of publications involving DL, GNNs, GNNs in the medical domain, overall multimodal and multimodal in biomedical and clinical sciences in the period 2015-2024 [18].
Multimodal Learning (MML) enhances task accuracy and reliability by leveraging information from various data sources or modalities [19]. This approach has witnessed a surge in popularity, as indicated by the growing body of MML-related publications (see Figure  1). By facilitating the fusion of multimodal data, such as radiological images, digitized pathology slides, molecular data, and electronic health records (EHR), MML offers a richer understanding of complex problems [20]. It enables the extraction and integration of relevant features that might be overlooked when analyzing data modalities separately. Recent advancements in MML, powered by Deep Neural Networks (DNNs), have shown remarkable capability in learning from diverse data sources, including computer vision (CV) and natural language processing (NLP) [21, 22]. Prominent multimodal foundation models such as Contrastive Language-Image Pretraining (CLIP) and Generative Pretraining Transformer (GPT-4) by OpenAI have set new benchmarks in the field [23, 21]. Additionally, the Foundational Language And Vision Alignment Model (FLAVA) represents another significant stride, merging vision and language representation learning to facilitate multimodal reasoning [24]. Within the realm of oncology, innovative applications of MML are emerging. The RadGenNets model, for instance, integrates clinical and genomics data with Positron Emission Tomography (PET) scans and gene mutation data, employing a combination of Convolutional Neural Networks (CNNs) and Dense Neural Networks to predict gene mutations in Non-small cell lung cancer (NSCLC) patients [25]. Moreover, GNNs and Transformers are being explored for a variety of oncology-related tasks, such as tumor classification [26], prognosis prediction [27], and assessing treatment response [28].
Recent literature has seen an influx of survey and review articles exploring MML [29, 30, 31, 32, 33]. These works have provided valuable insights into the evolving landscape of MML, charting key trends and challenges within the field. Despite this growing body of knowledge, there remains a notable gap in the literature regarding the application of advanced multimodal DL models, such as Graph Neural Networks (GNNs) and Transformers, in the domain of oncology. Our article aims to fill this gap by offering the following contributions:
@@
    Identifying large-scale MML approaches in oncology. We provide an overview of the state-of-the-art MML with a special focus on GNNs and Transformers for multimodal data fusion in oncology.
@@
    Highlighting the challenges and limitations of MML in oncology data fusion. We discuss the challenges and limitations of implementing multimodal data-fusion models in oncology, including the need for large datasets, the complexity of integrating diverse data types, data alignment, and missing data modalities and samples.
@@
    Providing a taxonomy for describing multimodal architectures. We present a comprehensive taxonomy for describing MML architectures, including both traditional ML and DL, to facilitate future research in this area.
@@
    Identifying future directions for multimodal data fusion in oncology. We identify GNNs and Transformers as potential solutions for comprehensive multimodal integration and present the associated challenges.
By addressing these aspects, our article seeks to advance the understanding of MML’s potential in oncology, paving the way for innovative solutions that could revolutionize cancer diagnosis and treatment through comprehensive data integration.
Our paper is organized as follows. Section 1 covers the fundamentals of MML, including data modalities, taxonomy, data fusion stages, and neural network architectures. Section 2 focuses on GNNs in MML, explaining graph data, learning on graphs, architectures, and applications to unimodal and multimodal oncology datasets. Section 3 discusses Transformers in MML, including architecture, multimodal Transformers, applications to oncology datasets, and methods of fusing data modalities. Section 4 highlights challenges in MML, such as data availability, alignment, generalization, missing data, explainability, and others. Section 5 provides information on data sources. Finally, we conclude by emphasizing the promise of integrating data across modalities and the need for scalable DL frameworks with desirable properties.
1 Fundamentals of Multimodal Learning (MML)
(a) An oerview of data collected from population to a tissue.
(b) Detailed look into data modalities acquired for cancer care.
Figure 2: We present various data modalities that capture specific aspects of cancer at different scales. For example, radiological images capture organ or sub-organ level abnormalities, while tissue analysis may provide changes in the cellular structure and morphology. On the other hand, various molecular data types may provide insights into genetic mutations and epigenetic changes.
@@ 1.1 Data Modalities in Oncology
A data modality represents the expression of an entity or a particular form of sensory perception, such as the characters’ visual actions, sounds of spoken dialogues, or the background music [34]. A collective notion of these modalities is called multi-modality [31]. Traditional data analysis and ML methods to study cancer data use single data modalities (e.g., EHR [35], radiology [12], pathology [36], or molecular, including genomics [37], transcriptomics [38], proteomics [39], etc.). However, the data is inherently multimodal, as it includes information from multiple sources or modalities that are related in many ways. Figure 2(a) provides a view of multiple modalities of cancer at various scales, from the population level to single-cell analysis. Oncology data can be broadly classified into 3 categories: clinical, molecular, and imaging, where each category provides complementary information about the patient’s disease. Figure 2(b) highlights different clinical, molecular, and imaging modalities. Multimodal analysis endeavors to gain holistic insights into the disease process using multimodal data.
@@ 1.1.1 Molecular Data
Molecular data modalities provide information about the underlying genetic changes and alterations in the cancer cells [40]. Efforts toward integrating molecular data resulted in the multi-omics research field [41]. Two principal areas of molecular analysis in oncology are proteomics and genomics. Proteomics is the study of proteins and their changes in response to cancer, and it provides information about the biological processes taking place in cancer cells. Genomics is the study of the entire genome of cancer cells, including changes in DNA sequence, gene expression, and structural variations [29]. Other molecular modalities include transcriptomics, pathomics, radiomics and their combinations, radiogenomics, and proteogenomics. Many publicly available datasets provide access to molecular data, including the Proteomics Data Commons for proteomics data and the Genome Data Commons for genetic data [42, 43].
@@ 1.1.2 Imaging Data
Imaging modalities play a crucial role in diagnosing and monitoring cancer. The imaging category can be divided into 2 main categories: (1) radiological imaging and (2) digitized histopathology slides, referred to as Whole Slide Imaging (WSI). Radiological imaging encompasses various techniques such as X-rays, CT scans, MRI, PET, and others, which provide information about the location and extent of cancer within the body. These images can be used to determine the size and shape of a tumor, monitor its growth, and assess the effectiveness of treatments. Histopathological imaging is the examination of tissue samples obtained through biopsy or surgery [44, 45]. Digitized slides, saved as WSIs, provide detailed information about the micro-structural changes in cancer cells and can be used to diagnose cancer and determine its subtype.
@@ 1.1.3 Clinical Data
Clinical data provides information about the patient’s medical history, physical examination, and laboratory results, saved in the patient’s electronic health records (EHR) at the clinic. EHR consists of digital records of a patient’s health information stored in a centralized database. These records provide a comprehensive view of a patient’s medical history, past diagnoses, treatments, laboratory test results, and other information, which helps clinicians understand the disease [46]. Within EHR, time-series data may refer to the clinical data recorded over time, such as repeated blood tests, lab values, or physical attributes. Such data informs the changes in the patient’s condition and monitors the disease progression [47].
1.2 Taxonomy of MML
We follow the taxonomy proposed by William et al. [34] (see Figure 3), which defines 5 main stages of multimodal classification: preprocessing, feature extraction, data fusion, primary learner, and final classifier, as given below:
@@ 1.2.1 Pre-processing
Pre-processing involves modifying the input data to a suitable format before feeding it into the model for training. It includes data cleaning, normalization, class balancing, and augmentation. Data cleaning removes unwanted noise or bias, errors, and missing data points [48]. Normalization scales the input data within a specific range to ensure that each modality contributes equally to the training [49]. Class balancing is done in cases where one class may have a significantly larger number of samples than another, resulting in a model bias toward the dominant class. Data augmentation artificially increases the size of the dataset by generating new samples based on the existing data to improve the model’s robustness and generalizability [48].
@@ 1.2.2 Feature Extraction
Different data modalities may have different features, and extracting relevant features may improve model learning. Several manual and automated feature engineering techniques generate representations (or embeddings) for each data modality. Feature engineering involves designing features relevant to the task and extracting them from the input data. This can be time-consuming but may allow the model to incorporate prior knowledge about the problem. Text encoding techniques, such as bag-of-words, word embeddings, and topic models [50, 51], transform textual data into a numerical representation, which can be used as input to an ML model [52]. In DL, feature extraction is learned automatically during model training[53].
@@ 1.2.3 Data Fusion
Data fusion combines raw features, extracted features, or class prediction vectors from multiple modalities to create a single data representation. Fusion enables the model to use the complementary information provided by each modality and improve its learning. Data fusion can be done using early, late, or intermediate fusion. Section 1.3 discusses these fusion stages. The choice of fusion technique depends on the characteristics of the data and the specific problem being addressed [54].
@@ 1.2.4 Primary Learner
The primary learner stage is training the model on the pre-processed data or extracted features. Depending on the problem and data, the primary learner can be implemented using various ML techniques. DNNs are a popular choice for primary learners in MML because they can automatically learn high-level representations from the input data and have demonstrated state-of-the-art performance in many applications. CNNs are often used for image and video data, while recurrent neural networks (RNNs) and Transformers are commonly used for text and sequential data. The primary learner can be implemented independently for each modality or shared between modalities, depending on the problem and data.
@@ 1.2.5 Final Classifier
The final stage of MML is the classifier, which produces category labels or class scores and can be trained on the output of the primary learner or the fused data. The final classifier can be implemented using a shallow neural network, a decision tree, or an ensemble model [34]. Ensemble methods, such as stacking or boosting, are often used to improve and robustify the performance of the final classifier. Stacking involves training multiple models and then combining their predictions at the output stage, while boosting involves repeatedly training weak learners and adjusting their weights based on the errors made by previous learners [55].
1.3 Data Fusion Strategies
@@
Fusion in MML can be performed at different levels, including early (feature level), intermediate (model level), or late (decision level) stages, as illustrated in figure 3. Each fusion stage has its advantages and challenges, and the choice of fusion stage depends on the characteristics of the data and the task.
@@
Figure 3: Taxonomy, stages, and techniques of multimodal data fusion are presented. Early, late, cross-modality fusion methods integrate individual data modalities (or extracted features) before, after, or at the primary learning step, respectively.
1.3.1 Early Fusion
@@
The early fusion involves merging features extracted from different data modalities into a single feature vector before model training. The feature vectors of the different modalities are combined into a single vector, which is used as the input to the ML model [34]. This approach can be used when the modalities have complementary information and can be easily aligned, such as combining visual and audio features in a video analysis application. The main challenge with early fusion is ensuring that the features extracted from different modalities are compatible and provide complementary information.
1.3.2 Intermediate Fusion